OpenStudy (anonymous):

I will award you! :) In the formula for average rate of change, what does the triangle in front of the x and y stand for? a. interval between x and y c. delta notation-change in y over change in x b. initial value of x and y d. the percent rate of change of y and x

OpenStudy (anonymous):

@SolomonZelman @lacrosseplayer22

OpenStudy (agreene):

Delta notation is used to indicate change: for instance, slope is described as \[\frac{\Delta y}{\Delta x}\] which means: \[\frac{y_2 - y_1}{x_2-x_1}\]

OpenStudy (agreene):

another example would be the distance formula: \[\sqrt{\Delta^2 x - \Delta^2 y} \rightarrow \sqrt{(x_2-x_1)^2-(y_2-y_1)^2}\]

OpenStudy (anonymous):

so it would be c?

OpenStudy (agreene):

indeed.